### v3

@@ 4.1 Efficiently performing FAGC
The FAGC module is combined with various state-of-the-art learning methods on the CIFAR-10 and CIFAR-100 datasets [45]. CIFAR-10 and CIFAR-100 are benchmark datasets comprising 50,000 training and 10,000 test images, which are in color and have a resolution of 32×32 pixels. CIFAR-100 shares the identical image set with CIFAR-10, consisting of 100 categories, with each category comprising 600 images. The training set is partitioned into labeled and unlabeled subsets, respectively. The numbers of samples per category for two experiments on the CIFAR-10 are 4 and 20, respectively, Thus, the total labeled samples are 40 and 200 for the two experiments, respectively. Similarly, for the CIFAR-100 dataset, we maintain the same sample ratios, yielding labeled samples of 400 and 2000 for the two scenarios.
We employ the ViT [44] network as the feature extraction model for CIFAR-10 and CIFAR-100 datasets, which are preprocessed with the methods of horizontal flips and random crops. Considering the image size of the CIFAR-100 dataset, the input scale of the standard ViT is adjusted to accommodate 32×32 image input and patch size of 2. The dimensions of the intermediate hidden layer, its depth, and the number of projection heads are set to 384, 12, and 6, respectively. The adjusted ViT is referred to as ViT-S [44]. Similarly, given the smaller number of categories in the CIFAR-10 dataset, we adopt the lightweight ViT-T [44] as the backbone network. It differs from the ViT-S structure in which the dimensions of the intermediate hidden layer are 192, and the number of projection heads is 3. We utilize the standard cross-entropy loss and utilize AdamW [61] as the optimizer. Additional configurations include a learning rate weight decay 0.0005, a learning rate 0.0005, a layer decay rate 0.5, and a moving average momentum 0.9.
@@
In our comparative analysis, the FAGC module is integrated with different learning models of which source codes are referenced from USB [62]. For each model, we gather experimental results from the last 10 epochs which are used to calculate the average accuracy. Simultaneously, we report a 95% confidence interval. The results are shown in Table 1. Obviously, FAGC combines different models to enhance the performance of the original model. Particularly, for the 400 labeled images in CIFAR-100, FAGC achieves an average classification accuracy improvement of approximately 4%. For the labeled images in CIFAR-10, FAGC exhibits a significant improvement when utilizing 40 labeled images. However, the enhancement of FAGC is not significant if the number of training samples is large, such as 200 labeled images in CIFAR-10. This is because the backbone network is more thoroughly trained with 400 images, and its accuracy reaches 90.13%.
@@
To describe the influence of the FAGC module combined with the neural networks, the FAGC module is combined with ViT-T, ViT-S, WRN-28-2 [63], and WRN-28-8 [63] in supervised training. As shown in Table 2, the performance of the models are significantly improved by the FAGC module. With the increase of the training samples, the influence of the FAGC module is decreased. Therefore, the FAGC module is more suitable for the small samples.
Table 1: Classification accuracy of the learning methods combined with the FAGC module on CIFAR-10 and CIFAR-100. "Supervised" denotes that only labeled data is used to train ViT-S and ViT-T, respectively. "AVG" describes the average accuracy.
@@
Table 2: Comparison of classification accuracy whether to use or not use the FAGC module.
@@
4.2 The effectiveness of FAGC in data augmentation methods
To evaluate the effectiveness of the FAGC module, we choose several representative augmentation methods which include horizontal flip, rotate, random crop, color jitter, cutout and augmix [64]. An experiment is executed by ViT-S with some data augmentation methods and FAGC for 400 labeled images in CIFAR-10. Table  3 describes the comparison of these methods whether to combine or not combine the FAGC module. From Table 3, the FAGC module can improve the classification accuracy combined with other data augmentation methods. It is noteworthy that the combination of all data augmentation methods does not achieve optimal classification accuracy, as shown in ID 8. In contrast, optimal accuracy is achieved with the combination of the methods of horizontal flip, random crop and the FAGC module.
Table 3: Comparison of the data augmentations whether to combine or not combine the FAGC module.
ID Data augmentations Methods
@@ 4.2 The effectiveness of FAGC in data augmentation methods
4.3 The influence of machine learning models on FAGC
Two experiments that choose 40 and 200 images from CIFAR-10 as training sets, respectively, are executed to understand the influence of the FAGC module on the feature classifiers. The feature extraction network model is ViT-S, and the feature classifiers for comparison include KNN [65], SVC [66], DecisionTree [67], ExtraTree [68], RandomForest [69], Bagging [70], GradientBoost [71] and MLP [72]. The results indicate that the FAGC module can improve the performance of the machine learning models, as shown in Table 4. Moreover, the MLP with the FAGC module as the highest classification accuracy. Since MLP uses the output layer weight parameters of the ViT-S.
Table 4: Comparison of different feature classifiers on CIFAR-10.
@@ 4.3 The influence of machine learning models on FAGC
4.4 The influence of the number of generated features on the model
@@
To evaluate the influence of the number of generated features for the models, we use multiple machine learning models as classifiers trained on 400 labeled features in CIFAR-100, including KNN, SVC, DecisionTree, ExtraTree, RandomForest, Bagging, GradientBoost and MLP. The feature extraction network model is ViT-S, and the numbers of generated features are 10, 100, 400, 1000 and 2000. The results of these experiments are presented in Figure 4. The generated feature quantity does not exhibit a strict correlation with the model’s classification accuracy. Excessive generated features can cause the model to overly focus on their distribution, but ignore the distribution of real features. Moreover, optimal performance in classification accuracy is achieved by MLP. Supplementary experiments are conducted utilizing different feature extraction models, including ViT-T and WRN-28-2. The detailed experimental results are shown in Figure 5. The results indicate that varying quantities of generated features are required by different methods, and for MLP the optimal number of generated features is 100.
Figure 4: The classification accuracy of the different numbers of the generated features.
Figure 5: The classification accuracy of the different numbers of the generated features for MLP.
4.5 Ablation study
@@
To evaluate the influence of the generated features from the Geodesic curves on the performance of models, experiments are conducted for the 40 labeled images on the CIFAR-10. ViT-T and ViT-S are selected as feature extraction models, and MLP is employed as the final classifier. Ablation experiments are conducted on the parameter $\lambda$ in the classification loss, as the Formula (10). It is noteworthy that if $\lambda=0$, the generated features are excluded from the classification training; on the other hand, if $\lambda=1$, the generated features are trained as real features. The experimental results are shown in Table  5. As the influence factor $\lambda$ increases, the classification accuracy of the model initially rises and then exhibits a slight decline. The declining trend is attributed to the increasing influence of the generated features on the loss function, causing the model to ignore the real features.
Table 5: The influence factor $\lambda$ on the classification performance on CIFAR-10.
Network
$\lambda$
0
0.1
0.3
0.45
0.5
0.55
0.7
0.9
1
ViT-T
79.32
80.95
81.86
82.92
82.83
82.85
82.63
82.46
82.32
ViT-S
90.00
92.34
93.56
93.68
93.77
93.79
93.54
93.23
93.11
@@ 5 Conclusion
In this paper, we propose a data augmentation method FAGC, which can provide additional information for the models of machine learning. Experimental results show that FAGC exhibits excellent performance on different datasets. In addition, the combination of the FAGC method with other data augmentation techniques can also achieve positive results. The enhanced generality and effectiveness can provide more comprehensive enhancement effect for machine learning tasks.
There are still potential improvements for the proposed method. For instance, the influence caused by isolated points. Alternatively, if image features are excessively scattered, and the Geodesic curve fails to reflect their distribution. It is more advisable to fit these points with a Geodesic surface. Improvements in these aspects will further boost the performance of FAGC and make it more applicable to various machine learning tasks.

### v4

@@ 4.2 Applying the FAGC in existing semi-supervised or supervised methods
The FAGC framework is combined with various state-of-the-art semi-supervised or supervised learning methods on the CIFAR-10 and CIFAR-100 datasets [46]. CIFAR-10 and CIFAR-100 are benchmark datasets comprising 50,000 training and 10,000 test images, which are in color and have a resolution of 32×32 pixels. CIFAR-100 shares the identical image set with CIFAR-10, consisting of 100 categories, with each category comprising 600 images. The training set is partitioned into labeled and unlabeled subsets, respectively. The numbers of samples per category for two experiments on the CIFAR-10 are 4 and 20, respectively, Thus, the total labeled samples are 40 and 200 for the two experiments, respectively. Similarly, for the CIFAR-100 datasets, we maintain the same sample ratios, yielding labeled samples of 400 and 2000 for the two scenarios. Considering the image size of the CIFAR dataset, the input scale of the standard ViT is adjusted to accommodate 32×32 image input and patch size of 2. We use ViT-T and ViT-S respectively to process CIFAR-10 and CIFAR-100. The semi-supervised methods include CoMatch [47], $\Pi$-Model [48], Mean Teacher [49], Pseudo Labeling [50], MixMatch [51], VAT [52], Dash [53], CRMatch [54], UDA [55], FixMatch [56], FlexMatch [57], FreeMatch [58], AdaMatch [59], SimMatch [60], SoftMatch [61].
@@
In our comparative analysis, the FAGC is integrated with different semi-supervised models of which source codes are referenced from USB [62]. For each model, we gather experimental results from the last 10 epochs which are used to calculate the average accuracy. Simultaneously, we report a 95% confidence interval. The results are shown in Table 1. Obviously, the FAGC can combine different models to enhance the performance of the original model. Particularly, for the 400 labeled images in CIFAR-100, the FAGC achieves an average classification accuracy improvement of approximately 4%. For the labeled images in CIFAR-10, the FAGC exhibits a significant improvement of about 1.5% when utilizing 40 labeled images. However, the enhancement of the FAGC is not significant if the number of training samples is large, such as 200 labeled images in CIFAR-10. This is because the backbone network is more thoroughly trained with 400 images, and its accuracy reaches 90.13%.
@@
To describe the influence of the FAGC combined with different backbone networks, the FAGC is combined with ViT-T, ViT-S, WRN-28-2 [63], and WRN-28-8 [63] in supervised training. As shown in Table 2, the performance of the models are significantly improved by the FAGC. With the increase of the training samples, the influence of the FAGC is decreased. Therefore, the FAGC delivers its maximum effectiveness under conditions of extreme data scarcity.
Table 1: The impact of the FAGC on the classification accuracies of different semi-supervised models. “AVG” (the average accuracy of all methods) is used to demonstrate the universal improvement effect achieved by adding the FAGC across different methods.
@@
Table 2: The impact of the FAGC on the classification accuracies of different backbone models.
@@
Table 3: Comparison of the FAGC with other feature augmentation methods on different datasets. “STL”, “ANI” and “CUB” represent datasets STL-10, Animals-90, and CUB-200-2011, respectively.
Methods STL ANI CUB
Baseline 94.73±0.05 96.40±0.09 67.49±0.31
Mani-Mix-up [33] 94.39±0.08 96.60±0.15 68.17±0.25
SFA-S [30] 94.96±0.05 97.20±0.16 68.85±0.19
MoEx [31] 94.44±0.18 96.60±0.07 67.95±0.21
FL-SMOTE [32] 94.60±0.12 96.80±0.08 68.17±0.23
FAGC 95.80±0.15 97.00±0.07 69.30±0.28
4.3 Comparison with other feature augmentation methods on additional datasets
To further validate the effectiveness and applicability of the FAGC, comparisons with other feature augmentation methods are conducted on several more challenging datasets. The datasets include STL-10 (STL) [64], CUB-200-2011 (CUB) [65], and Animals-90 (ANI) [66]. STL-10 contains 10 categories of data, with 500 images for training and 800 images for testing in each category. CUB-200-2011 contains a total of 11788 images belonging to 200 categories of birds, of which 5994 are used for training and the rest for testing. Animals-90 is a collection of 5400 animal images from 90 categories collected by google images. The feature augmentation methods used for comparison include Mani-Mix-up [33], SFA-S [30], MoEx [31], and FL-SMOTE [32].
We extracted a subset from each dataset for training to demonstrate the effectiveness of the FAGC in extremely small-sample scenarios. The subset consists of 10 categories, with only 4 labeled samples per category for training. The feature extraction network used is ViT-S, and the classifier is the MLP. The results are shown in Table 3, the addition of the FAGC further improved the classification accuracy when facing more challenging data. Regardless of the initial accuracy, the improvement of model performance by the FAGC demonstrates its applicability. When compared with various feature augmentation methods, the FAGC achieved the most significant performance improvement in the majority of cases.
4.4 The combination of the FAGC and common pixel-level augmentation methods.
As a feature-level augmentation method, the FAGC can be combined with existing pixel-level augmentation methods. We choose several representative pixel-level augmentation methods which include horizontal flip, rotate, random crop, color jitter, cutout and augmix [67]. The experiments deploy ViT-S network on the CIFAR-10 dataset with 400 labeled data. Table  4 describes the comparison of performance before and after combining these methods with the FAGC. From Table  4, the FAGC can improve the classification accuracy combined with other pixel-level augmentation methods. It is noteworthy that the combination of all pixel-level augmentation methods does not achieve optimal classification accuracy, as shown in ID 8 of Table  4. In contrast, optimal accuracy is achieved with the combination of the methods of horizontal flip, random crop and the FAGC.
Table 4: Comparison of the compatibility between the FAGC and various pixel-level augmentation methods.
ID Pixel-level augmentation methods Methods
@@ 4.3 Comparison with other feature augmentation methods on additional datasets
4.5 Different machine learning classifiers with the FAGC
[…]
Table 5: The impact of the different feature classifiers on the performance with the FAGC on CIFAR-10.
@@ 4.5 Different machine learning classifiers with the FAGC
4.6 The different number of the augmented feature vectors
@@
To evaluate the influence of the number of augmented feature vectors for the models, we use multiple machine learning models as classifiers trained on 400 labeled features in CIFAR-100, including KNN, SVC, DecisionTree, ExtraTree, RandomForest, Bagging, GradientBoost and MLP. The feature extraction network model is ViT-S, and the numbers of the augmented feature vectors are 10, 100, 400, 1000, and 2000. The results of the experiments are presented in Figure 4. During the process of increasing the number from 10 to 2000, the performance of all classifiers exhibited a consistent pattern: rapid initial improvement, followed by a gradual decline, and eventual stabilization. The number of the generated feature vectors does not exhibit a strict correlation with the model’s classification accuracy. Excessive augmented feature vectors can cause the model to overly focus on their distribution, but ignore the distribution of real feature vectors. Moreover, the best performance is achieved when using the MLP classifier with the number of augmented feature vectors set to 100. Supplementary experiments are conducted utilizing different feature extraction models, including ViT-T, ViT-S, WRN-28-8, and WRN-28-2. The number of the augmented feature vectors is also increased from 10 to 2000, while the classifier is consistently maintained as the MLP to observe the impact of different feature extraction models on the FAGC. The detailed experimental results are shown in Figure 5. The results indicate that the performance improvement of the FAGC varies across different feature extraction models, with the most significant enhancement observed for ViT-T. As the number of the augmented feature vectors increases, the gain brought by the FAGC aligns with the conclusions from Figure 4, achieving optimal effectiveness when the number is set to 100.
Figure 4: The impact of the number of the augmented feature vectors on model performance under different feature classifiers. The X-axis represents the number of the augmented feature vectors, and the Y-axis represents the classification accuracy with different classifiers.
Figure 5: The impact of the number of augmented features on model performance under different feature extraction models. The X-axis represents the number of the augmented feature vectors, and the Y-axis represents the classification accuracy with different feature extraction models.
4.7 The influence of factor $\lambda$
@@
To evaluate the influence of the augmented features from the Geodesic curves on the performance of models, the experiments are conducted for the 40 labeled images on the CIFAR-10. ViT-T and ViT-S are selected as feature extraction models, and MLP is employed as the final classifier. The ablation experiments are conducted on the parameter $\lambda$ in the classification loss, as the Formula (10). As the value of $\lambda$ increases, the model assigns greater weight to the knowledge derived from the augmented feature vectors. The experimental results are shown in Table  6. As the influence factor $\lambda$ increases, the classification accuracy of the model initially rises and then exhibits a slight decline. The declining trend is attributed to the increasing influence of the augmented features on the loss function, causing the model to ignore the real features.
Table 6: The influence of factor $\lambda$ on the classification performance on CIFAR-10.
Methods $\lambda$
0 0.1 0.3 0.45 0.5 0.55 0.7 0.9 1
ViT-T 79.32 80.95 81.86 82.92 82.83 82.85 82.63 82.46 82.32
ViT-S 90.00 92.34 93.56 93.68 93.77 93.79 93.54 93.23 93.11
Figure 6: The t-SNE visualization of the constructed Geodesic curve. The blue dots are the initial feature vectors, and the yellow dots are the augmented feature vectors sampled on the Geodesic curve.
Figure 7: Visualizations of the augmented feature vectors with the FAGC. (a) is the heatmaps generated directly from the feature vectors, and (b) is the reconstructed images obtained by decoding latent vectors using a pre-trained autoencoder.
4.8 Visualization
For a clearer comparison between augmented and initial feature vectors, t-SNE, heatmaps, and autoencoder (AE) are used to visualize the augmented feature vectors. Specifically, the t-SNE technique visualizes feature vectors extracted from the CIFAR dataset by a pre-trained ResNet-18. In Figure 6, there are 50 initial feature vectors and 150 augmented feature vectors. It can be observed that the constructed Geodesic curve, i.e., the path formed by the yellow dots, fits the initial feature vectors well and shows no sign of overfitting. Furthermore, Figure 7 reveals the semantic relationships between the augmented feature vectors and the initial feature vectors. The heatmaps in 7(a) are generated from the feature vectors of randomly selected cat images. In 7(b), the initial images are originated from the Oxford-IIIT Pets [76] dataset. The reconstructed images are generated by encoding the initial images using the pre-trained encoder, augmenting the feature vectors through the FAGC, and then reconstructing the images via the decoder. As demonstrated in Figure 7, the feature vectors augmented by the FAGC maintain semantic consistency with the original feature vectors, and their variation remains within controlled bounds. The points along the Geodesic curve, i.e., the augmented feature vectors, capture rich geometric transformation information between initial feature vectors with high confidence. Collectively, these visualization experiments validate the safety and reliability of the feature vectors augmented by the FAGC, demonstrating their suitability for many practical scenarios.
@@ 5 Conclusion
This paper proposes a feature augmentation method in the pre-shape space, called the FAGC, which provides additional information for machine learning and deep learning models. The FAGC encompasses the entire workflow of the feature extraction, shape modeling, Geodesic curve construction, and downstream tasks application, enabling its convenient deployment in diverse small-sample scenarios. Crucially, The Geodesic curve construction is the core innovation of the FAGC. Unlike existing approaches that simply connect pairs of data in the pre-shape space, the FAGC proposes a novel method to construct the Geodesic curves across multiple initial pre-shapes. The FAGC has been validated across numerous designed small-sample scenarios, with results demonstrating the effectiveness of the features augmented by the FAGC. Furthermore, visualization results across multiple scenarios further verify the safety and reliability of the FAGC.
The FAGC has successfully bridged the shape space theory with deep learning and machine learning, opening up new research avenues for data augmentation. However, the FAGC still has some areas for improvement. For instance, the Geodesic curves may not be the optimal geometric structure for data augmentation, and there is a need to explore diversified applications of the FAGC in deep learning models. In the future, we will explore more complex geometric structures in the pre-shape space, such as Geodesic surfaces, and investigate modular applications of the FAGC in deep learning frameworks.
6 acknowledgements
This research is sponsored by National Natural Science Foundation of China (Grant No. 52273228), Key Program of Science and Technology of Yunnan Province (202302AB080022), the Project of Key Laboratory of Silicate Cultural Relics Conservation (Shanghai University), Ministry of Education (No. SCRC2023ZZ07TS).